A small leak will fill an empty boat in 5 hours. The boat's bilge pump will empty a full boat in 3 hours. If the boat were half full and leaking, how long would it take the pump to empty the boat?
Answer by stanbon(75887)

A small leak will fill an empty boat in 5 hours. The boat's bilge pump will empty a full boat in 3 hours. If the boat were half full and leaking, how long would it take the pump to empty the boat?
---------------------------------------------
Fill rate = 1/5 job/hr
---
Empty rate = 1/3 job/hr
---
Together rate = 1/5 - 1/3 = (3-5)/15 = -2/15 job/hr.
---
Time to empty a full boat = 15/2 = 7 1/2 hrs
-----
Time to empty (1/2) boat = 15/4 = 3 3/4 hrs = 3 hr 45 minutes
======================
Cheers,
Stan H.